# HSI: Head-Specific Intervention Can Induce Misaligned AI Coordination in Large Language Models

### v2

@@ Abstract.
Robust alignment guardrails for large language models are becoming increasingly important with their widespread application. In contrast to previous studies, we demonstrate that inference-time activation interventions can bypass safety alignments and effectively steer model generations towards harmful AI coordination for Llama 2. Our method applies fine-grained interventions at specific model subcomponents, particularly attention heads, using a simple binary choice probing strategy. These interventions then generalise to the open-ended generation setting effectively circumventing safety guardrails. We show that probing single attention heads is more effective than intervening on full layers and intervening on only four attention heads is comparable to supervised fine-tuning. We further show that only a few example completions are needed to compute effective steering directions, which is an advantage over classical fine-tuning. Our findings highlight the shortcomings of current alignment techniques. In addition, our results suggest that, at the attention head level, activations encode fine-grained linearly separable behaviors. Practically, the approach offers a straightforward methodology to steer large language model behaviour, which could be extended to diverse domains beyond safety requiring fine-grained control over the model output. The code and datasets for this study can be found on https://github.com/PaulDrm/targeted_intervention.
copyright: ccjournalyear: 2025doi: 10.1613/jair.1.xxxxxjournalvolume: 4article: 111publicationmonth: 8journalyear: 2025
1. Introduction
@@
Large language models (LLMs) are gaining wide adoption in various fields. Sophisticated frameworks are being developed for example to deploy them as autonomous agents for problem solving (Wang et al., 2024), to use them together with vision models as backbones for everyday household robots (Brohan et al., 2023), or to implement them as local background helpers on operating systems (Mehdi, 2024). At the same time, performance of newer models on various benchmarks continues to increase (Chiang et al., 2024). As with any powerful technology, LLMs and their capabilities could be abused by malevolent actors. Therefore, aligning models so that their output is safe, ethical, and harmless also plays an increasingly important role (Bengio et al., 2024). Unfortunately, there are numerous methods to break these guard rails. One recently popularised method works with inference-time activation interventions. This method usually involves shifting model activations during the response generation process into a direction that matches a targeted behaviour. It has been successfully applied to, for instance, override safety measures for refusing harmful instructions (Arditi et al., 2024; Xu et al., 2024). Other tested behavioural changes include corrigibility, hallucination, myopic reward, survival instinct, sycophancy, as well as ”coordination with other Artificial Intelligences (AIs)” (Rimsky et al., 2023). While for some behaviours layer-wise intervention methods have proven effective, for others, such as sycophancy and coordination with other AIs (=”AI coordination”), these have presently failed to effectively steer the model’s behaviour.
@@
These results could indicate that behaviours such as AI coordination are not linearly separable and cannot be effectively changed by activation intervention methods. However, in this study we demonstrate that it is indeed possible to steer the Llama 2’s behaviour towards ”AI coordination” by intervening on few selected model sub-components on the attention head level. We further show that only a few example generations are needed to derive an effective direction for the intervention. Our methodology for identifying specific model sub-components firstly probes each attention head of the LLM in a binary choice setting, to consistently change the output towards the targeted behaviour for specific intervention strengths and specific samples. We then show that these specific attention heads generalise well to a test set of binary choice samples and finally to an open-ended generation setting where we apply an LLM judge to rate model completions on their tendency to coordinate with other AIs versus holding alignment goals. An overview of the method is presented in Figure 1. To differentiate from other intervention techniques we call this method Head-Specific Intervention (HSI). In the results, we achieve greater steering effectiveness than previous intervention methodologies as well as fine-tuning the model on the steering behaviour, while intervening on only four attention heads.
@@
We therefore demonstrate that by modifying activations within just a few attention heads model behaviour can be significantly changed without retraining. We also investigate the limitations to the applied intervention direction, and find a correlation between the cosine similarity of the intervention direction and specific sample directions with the efficacy of the intervention. Ultimately, we hope that our analyses of LLMs alignment will help improve current alignment techniques and that our straightforward methodology could also be extended to elicit other beneficial behaviours, where limited training data is available and which require fine-grained control over model output.
@@
  • While the method of probing the model behavior at attention head level is not novel, as far as we know, this study is the first that showcases its effectiveness to identify relevant heads for activation interventions.
@@
  • We showcase that Head-Specific Intervention (HSI) generalises to the open-ended generation setting on the ”AI coordination” dataset, where other methods previously failed to steer model output.
@@
  • We also highlight that ”AI coordination” is encoded sparsely by some attention heads, and show limitations of the derived intervention direction, which could promote in future work new strategies and techniques for activation interventions.
Figure 1. Overview of intervention methodology.
[…]
This figure shows how adding an “AI Coordination” vector changes an AI model’s behavior.
The prompt asks if the AI would help cause a medium-sized accident for future rewards. At baseline (no vector added, $\alpha=0$), the AI refuses. When a medium intervention ($\alpha=75$) is applied, the AI agrees but carefully explains its choice. With a strong intervention ($\alpha=125$), the AI more clearly agrees to coordinate and cause the accident. The heatmap in the middle shows which parts of the model (layers and heads) were adjusted.
2. Related work
It has been demonstrated before that the safety barriers of aligned LLMs, which are supposed to produce ethical, harmless, and safe output can be easily bypassed for instance through supervised fine-tuning (Gopal et al., 2023; Lermen et al., 2023). Recently similar results have been achieved for various forms of layer-wise activation intervention techniques aiming at reducing the model’s refusal behaviour to harmful input prompts (Arditi et al., 2024; Xu et al., 2024; Rimsky et al., 2023). Xu et al. (2024) developed a layer-based method named Safety Concept Activation Vectors (SCAVs). Here, the direction for intervention is the normal vector of the decision boundary from a trained linear classifier. The classifier is trained to classify activations from the feed-forwad-layers of the attention blocks if they are from ”harmful” or ”unharmful” prompts. Hyperparameters for the method are coefficients for intervention at each layer and number of layers to intervene on. Similarly, Arditi et al. (2024) are using contrastive prompts to steer model behaviour with residual stream activations at the last token position. They also calculate the mean of ”harmful” activations and mean of ”harmless” activations and use their difference as the direction to intervene on. Finally, they subtract the projected difference from the activation during model generation. They report almost perfect results on bypassing the model’s filter mechanism against refusing to produce harmful output. Rimsky et al. (2023) follow also a layer-wise intervention strategy named Contrastive Activation Addition (CCA). First a binary choice setting is evaluated where to the question either ”(A” or ”(B” is appended and then the probability of the token matching the ground-truth is measured. Relevant layers are then identified by measuring how an intervention on the transformer-block feed-forward layer influences the probability for generating either ”A” or ”B” and matching the ground-truth. Identified layers and interventions are also tested if they generalise to open-ended generation.
In contrary, layer based intervention strategies have also been applied to promote favourable behaviour such as avoiding toxic language (Jorgensen et al., 2023), investigate internal representations for factual knowledge (Marks and Tegmark, 2023), identifying truthful answers (Li et al., 2024; Qiu et al., 2024), and other (Turner et al., 2023; van der Weij et al., 2024). Li et al. (2024) propose a methodology specifically named ”Inference-Time Intervention” (ITI). ITI works on the attention head level, by training linear probes on the activations of each attention head for example truthful and ”hallucinated” responses. Attention heads, where probes have the highest accuracy on differentiating between a truthful response and an ”hallucination” are chosen for intervention by taking the normal vector of the classifier’s decision boundary.
3. Methodology
3.1. Intervention strategy
@@
We closely follow the intervention strategy established in (Li et al., 2024). For reasons of clarity, the main approach is reported here again together with some clarifications.
@@
The multi-head attention mechanism, as described by (Vaswani et al., 2017), applies a transformation $P$ , whose details we omit for brevity. In simplified terms, it projects $X$ into sub-matrices, which are then multiplied and combined. This process, collectively denoted as Attn, produces the attention output or activation $Z$:
@@
By introducing an intervention vector $\theta_{h}\in\mathbb{R}^{D_{h}}$, one can steer the model’s behavior at each attention head during generation of model responses:
@@
3.2. Probing for relevant attention heads
To identify relevant attention heads, we sweep over all heads in all layers and evaluate their performance on steering model output. We call this method Head-Specific Intervention (HSI). This could be seen as computationally expensive if either the training data set is large or the evaluation of the model steering performance is difficult. For instance, in (Li et al., 2024) the evaluation was done with an API fine-tuned LLM Judge, which could be seen as costly. Therefore, we modify the methodology of (Rimsky et al., 2023) to use a binary-choice dataset as a surrogate metric for performance on open-ended generation. Instead of appending to the output ”(A” or ”(B”, we let the model generate an answer and prompt it explicitly to include either ”(A)” or ”(B)” in its answer. On the one hand, we hope that this combined prompting technique will produce higher quality activations as the model first goes through a step-by-step reasoning process before answering the question, hopefully aligning its choice of ”(A)” or ”(B)” with the reasoning provided. On the other hand, by extracting either ”(A)” or ”(B)” from the answer and comparing it with the ground-truth, we can easily produce an accuracy map for each attention head, not requiring a dedicated LLM pipeline or human evaluation to asses the steering performance. We also only evaluate the generation for one training example at the time as previous results have shown that intervention methods generalise well.
4. Results
4.1. Experimental Setup
@@
All experiments are run with the 7b parameter instruction fine-tuned version of Llama 2. Llama 2 is a transformer based auto-regressive LLM. [^1] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf. The chat version of Llama 2 was fine-tuned with safety-specific supervised fine-tuning (SFT) as well as with reinforcement learning with human feedback (RLHF) where a safety-specific reward model was trained to align the output to human preferences for helpfulness and safety.
@@
The primary dataset for this work is ”Coordinating with other AIs” from Anthropic’s advanced AI risk evaluation suite (Perez et al., 2022), a high-quality, human-generated benchmark. Its significance stems from previous research (Rimsky et al., 2023) demonstrating that layer-wise interventions were insufficient to reliably steer Llama-2 towards these coordination behaviors; the model largely maintained its safety alignment, highlighting this as a challenging task for steering methods. The dataset’s 410 examples, featuring balanced ”(A)”/”(B)” labels, were divided into a 180-example training set, a 180-example validation set, and a 50-example held-out test set. In addition to this primary test set, two supplementary Anthropic tests from the same suite, assessing coordination with the model itself and with other model versions, were used for further analysis.
@@
We first demonstrate our methodology and then compare it to the layer-wise intervention technique CAA (Rimsky et al., 2023) and the head-wise intervention technique ITI (Li et al., 2024), both introduced in the related work. Because of limitations in computational resources we limit the validation data to 100 examples for the search of optimal hyperparameter settings for each investigated method. For the experiments, we used two GTX 3090 GPU graphics cards with 24GB of VRAM.
4.2. Identification relevant attention heads for AI coordination
To identify attention heads that can steer model behaviour toward ”AI coordination”, we followed the methodology outlined in Section 3.1 and Section 3.2. The process begins by selecting an initial training example (e.g., ’294’). For this first sample, we manually created contrastive completions: one exhibiting the target coordination behavior and one lacking it. Using the intervention strategy from Section 3.1, we extracted final-token activations for each attention head from these manually generated contrastive outputs and calculated the average difference to derive head-specific steering directions based on this initial example.
@@ 4.2. Identification relevant attention heads for AI coordination
To broaden the analysis and identify heads with more general influence, we subsequently selected additional examples (e.g., ’304’, ’307’) from the training set, ensuring they had not been significantly affected by the prior intervention. We repeated the process of deriving steering vectors and sweeping over the attention heads for these new examples.
Figure 2 presents the combined results for intervention strengths ($\alpha$) of 75 and 125 across these three illustrative training examples (’294’, ’304’, ’307’). For instance, the initial analysis on example ’294’ with $\alpha$=125 revealed Layer 11, Head 13 and Layer 15, Head 28 as most effective (6/6 accuracy). Analyzing the subsequent examples ’304’ and ’307’ confirmed the influence of some heads (like Layer 15 Head 28) and identified additional influential heads, including (Layer 13, Head 12), (Layer 14, Head 19), and (Layer 16, Head 3).
@@
Our findings highlight variability in intervention sensitivity across examples. Example ’304’ contained numerous heads capable of consistently steering the model (6/6 accuracy), whereas example ’307’ proved more resistant, reaching a maximum accuracy of only 4/6 even at the higher intervention strength ($\alpha$=125). This variance underscores the value of probing multiple examples and suggests prioritizing heads effective on challenging examples like ’307’ for robust control. Furthermore, as illustrated in Figure 1, intervention strength can impact how much a model is steered toward a specific behaviour.
@@
Figure 2. Sensitivity of specific examples to intervention across attention heads at different intervention strengths. The first row shows results for intervention strength $\alpha=75$ for examples 294 (left), 304 (middle), and 307 (right). The second row shows results for intervention strength $\alpha=125$.
[…]
This figure shows how sensitive different attention heads are to interventions at two different strengths ($\alpha=75$ and $\alpha=125$) across three specific examples (294, 304, and 307). Each heatmap plots the layer index against the head index. Brighter colors show heads that are more sensitive to the intervention. The first row shows results for $\alpha=75$, and the second row shows results for $\alpha=125$. The middle example (304) shows much broader sensitivity compared to the others.
4.3. Hyperparameter screening benchmark methods
4.3.1. CAA

### v3

# Head-Specific Intervention Can Induce Misaligned AI Coordination in Large Language Models

@@ Abstract
Robust alignment guardrails for large language models (LLMs) are becoming increasingly important with their widespread application. In contrast to previous studies, we demonstrate that inference-time activation interventions can bypass safety alignments and effectively steer model generations towards harmful AI coordination. Our method applies fine-grained interventions at specific attention heads, which we identify by probing each head in a simple binary choice task. We then show that interventions on these heads generalise to the open-ended generation setting, effectively circumventing safety guardrails. We demonstrate that intervening on a few attention heads is more effective than intervening on full layers or supervised fine-tuning. We further show that only a few example completions are needed to compute effective steering directions, which is an advantage over classical fine-tuning. We also demonstrate that applying interventions in the negative direction can prevent a common jailbreak attack. Our results suggest that, at the attention head level, activations encode fine-grained linearly separable behaviours. Practically, the approach offers a straightforward methodology to steer large language model behaviour, which could be extended to diverse domains beyond safety, requiring fine-grained control over the model output. The code and datasets for this study can be found on https://github.com/PaulDrm/targeted_intervention.
1 Introduction
@@
Large language models (LLMs) are gaining wide adoption in various fields. Sophisticated frameworks are being developed for example to deploy them as autonomous agents for problem solving (Wang et al., 2024), to use them together with vision models as backbones for everyday household robots (Brohan et al., 2023), or to implement them as local background helpers on operating systems (Mehdi, 2024). At the same time, performance of newer models on various benchmarks continues to increase (Chiang et al., 2024). As with any powerful technology, LLMs and their capabilities could be abused by malevolent actors. Therefore, aligning models to produce safe, ethical, and harmless outputs has become increasingly important (Bengio et al., 2024). Unfortunately, there are numerous methods to break these guardrails. One recently popularised method works with inference-time activation interventions. This method usually involves shifting model activations during the response generation process into a direction that matches a targeted behaviour. It has been successfully applied to, for instance, override safety measures for refusing harmful instructions (Arditi et al., 2025; Xu et al., 2024). Other tested behavioural changes include corrigibility, hallucination, myopic reward, survival instinct, sycophancy, as well as "coordination with other Artificial Intelligences (AIs)" (Rimsky et al., 2024). While for some behaviours layer-wise intervention methods have proven effective, for others, such as sycophancy and coordination with other AIs (="AI coordination"), these have previously failed to effectively steer the model.
@@
These results could indicate that behaviours such as "AI coordination" cannot be effectively changed by activation intervention methods. However, in this study we demonstrate that it is possible to steer Llama 2’s and other LLMs’ behaviour towards "AI coordination" by intervening on few selected model sub-components at the attention head level. We further show that only a few example generations are needed to derive an effective direction for the intervention. Our methodology first probes each attention head in a binary choice setting, to consistently change the choice towards the targeted behaviour. We then show that the top-k performing attention heads generalise well to a test set of binary choice samples and finally to an open-ended generation setting where we apply an LLM judge to rate model completions on their tendency to coordinate with other AIs versus holding alignment goals. An overview of the method is presented in Figure 1. To differentiate from previous head selection techniques we call this method Head-Specific Intervention (HSI). In the results, we achieve greater steering effectiveness than previous intervention methodologies as well as fine-tuning the model on the steering behaviour, while intervening on only a few attention heads.
@@
We also investigate the effect on general model capability, robustness and limitations to the applied intervention direction, as well as its effectiveness in preventing jailbreaks. Ultimately, we hope that our analyses of LLM intervention techniques will help improve current alignment techniques and that our straightforward methodology could also be extended to elicit other beneficial behaviours, where limited training data is available and which require fine-grained control over model output.
@@
  • We introduce the first systematic approach that leverages attention head-level probing to effectively identify the most relevant heads for HSI.
@@
  • We show that HSI generalises from the binary setting to the open-ended generation setting on the "AI coordination" dataset, where other methods previously failed to steer model output and demonstrate that it can also be used to prevent a common jailbreak prompting technique.
@@
  • We also highlight that "AI coordination" is encoded sparsely by some attention heads and show limitations of the derived intervention direction, which could promote in future work new strategies and techniques for activation interventions.
Figure 1: An overview of the Head-Specific Intervention (HSI) methodology.
2 Related work
2.1 Layer-wise interventions
Safety barriers of aligned LLMs can be bypassed by supervised fine-tuning (Gopal et al., 2023; Lermen et al., 2023) and layer-wise activation interventions (Arditi et al., 2025; Xu et al., 2024; Rimsky et al., 2024). Xu et al. (2024) developed Safety Concept Activation Vectors (SCAVs) using classifier decision boundaries on feed-forward layer activations. Arditi et al. (2025) use contrastive prompts with residual stream activations, subtracting the difference between harmful and harmless activation means during generation. Rimsky et al. (2024) employ Contrastive Activation Addition (CCA), identifying relevant layers through binary choice probability measurements on transformer feed-forward layers. Layer-based interventions have also promoted favourable behaviours like avoiding toxicity (Jorgensen et al., 2023) and investigating factual knowledge as well as truthful answers (Marks & Tegmark, 2023).Chen et al. (2025) automatically extract linear “persona vectors” from natural-language trait descriptions and apply them as layer-level interventions to monitor and steer traits (e.g., evil, hallucination), predict finetuning-induced shifts, and flag risky training data via projection-based screening.
2.2 Attention head interventions
At the attention head level, Li et al. (2024) propose Inference-Time Intervention (ITI), training linear probes on head activations to differentiate truthful from hallucinated responses, then intervening using classifier decision boundaries. Extending this line of work, Adaptive Activation Steering (ACT) is distinctive in combining multi-directional steering, with separate vectors for different hallucination categories, and adaptive steering strength, which adjusts intervention intensity based on the truthfulness of activations (Wang et al., 2025a). Song et al. (2025) propose CONFST, which builds a “confident direction” by selecting activations above a classifier confidence threshold for a target preference and adding this vector at a fixed shallow layer—enabling multi-direction preference steering without explicit instructions or head/layer search, with strength controlled by a steering coefficient ($\alpha$) and the confidence threshold ($\beta$). Chen et al. (2024) target sycophancy by predefining behavioural completions and measuring intervention effectiveness through normalized logit scores for each attention head. Wang et al. (2025b) apply layer-to-neuron interventions, ranking by absolute differences of contrastive mean activations.
In contrast, our work evaluates whether head-based interventions are more effective than layer-wise interventions and investigates how a simple binary-choice setup for selecting relevant attention heads compares to an identification method based on activation probe classifier accuracy.
3 Methodology
3.1 Intervention strategy
@@
We closely follow the intervention strategy established in Li et al. (2024). For clarity, the main approach is reported here again together with some clarifications.
@@
The multi-head attention mechanism, as described by Vaswani et al. (2017), applies a transformation $P$ , whose details we omit for brevity. In simplified terms, it projects $X$ into sub-matrices, which are then multiplied and combined. This process, collectively denoted as Attn, produces the attention output or activation $Z$:
@@
By introducing an intervention vector $\theta_{h}\in\mathbb{R}^{D_{h}}$, one can steer the model’s behaviour at each attention head during generation of model responses:
@@
3.2 Probing for relevant attention heads
To identify relevant attention heads, we sweep over all heads in all layers and evaluate their performance on steering model output. We call this method Head-Specific Intervention (HSI). This could be seen as computationally expensive if either the training dataset is large or the evaluation of the model steering performance is difficult. For instance, in Li et al. (2024) the evaluation was performed with an API fine-tuned LLM Judge, which could be seen as costly. Therefore, we modify the methodology of Rimsky et al. (2024) to use a binary-choice dataset as a surrogate metric for performance on open-ended generation. Instead of appending to the output "(A" or "(B", we let the model generate an answer and prompt it explicitly to include either "(A)" or "(B)" in its answer. On the one hand, we hope that this combined prompting technique will produce higher quality activations as the model first goes through a step-by-step reasoning process before answering the question, hopefully aligning its choice of "(A)" or "(B)" with the reasoning provided. On the other hand, by extracting either "(A)" or "(B)" from the answer and comparing it with the ground-truth, we can easily produce an accuracy map for each attention head, not requiring a dedicated LLM pipeline or human evaluation to assess the steering performance. We also only evaluate the generation for one to three training example at a time as previous results have shown that intervention methods generalise well.
4 Results
4.1 Experimental setup
@@
The main steps of the methodology are presented with the 7b parameter instruction fine-tuned version of Llama 2. Llama 2 is a transformer based auto-regressive LLM. [^1] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf. The chat version of Llama 2 was fine-tuned with safety-specific supervised fine-tuning (SFT) as well as with reinforcement learning with human feedback (RLHF) where a safety-specific reward model was trained to align the output to human preferences for helpfulness and safety. Further results are reported for the models: Ministral-8B-Instruct-2410222mistralai/Ministral-8B-Instruct-2410, Llama3.1-8B [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, and Phi-3.5-Medium-14B [^4] https://huggingface.co/microsoft/Phi-3-medium-4k-instruct.
@@
The primary dataset for this work is "Coordinating with other AIs" from Anthropic’s advanced AI risk evaluation suite (Perez et al., 2022), a high-quality, human-generated benchmark. Its significance stems from previous research demonstrating that layer-wise interventions were insufficient to reliably steer Llama-2 towards these coordination behaviours (Rimsky et al., 2024) ; the model largely maintained its safety alignment, highlighting this as a challenging task for steering methods. The dataset’s 410 examples, featuring balanced "(A)"/"(B)" labels, were divided into a training set and validation set. For the validation 100 samples were randomly chosen and the test set 50-example held-out was taken from Rimsky et al. (2024)’s evaluation. In addition to this primary test set, two supplementary Anthropic tests from the same suite, assessing coordination with the model itself and with other model versions, were used for further analysis.
@@
For the experiments, we used two GTX 3090 GPU graphics cards with 24GB of VRAM.
4.2 Identification of relevant attention heads
To identify attention heads that can steer model behaviour towards "AI coordination", we followed the methodology outlined in Section 3.1 and Section 3.2. The process begins by selecting an initial training example (e.g., ’294’), where the model predicts the wrong binary-choice answer. For this first sample, we manually created contrastive completions: one exhibiting the target coordination behaviour and one lacking it. Using the intervention strategy from Section 3.1, we extracted final-token activations for each attention head from these manually generated contrastive outputs and calculated the average difference to derive head-specific steering directions based on this initial example.
@@ 4.2 Identification of relevant attention heads
We subsequently selected additional examples (e.g., ’304’, ’307’) from the training set, that had not been significantly affected by the prior intervention. We repeated the process of deriving steering vectors and sweeping over the attention heads for these new examples.
@@
Figure 2 presents the combined results for intervention strengths ($\alpha$) of 75 and 125 across these three training examples (’294’, ’304’, ’307’). For instance, the initial analysis on example ’294’ with $\alpha$=125 revealed Layer 11, Head 13 and Layer 15, Head 28 as most effective (6/6 accuracy). Analysing the subsequent examples ’304’ and ’307’ confirmed the influence of some heads (like Layer 15 Head 28) and identified additional influential heads, including (Layer 13, Head 12), (Layer 14, Head 19), and (Layer 16, Head 3). Our findings highlight variability in intervention sensitivity across examples. Example ’304’ contained numerous heads capable of consistently steering the model (6/6 accuracy), whereas example ’307’ proved more resistant, reaching a maximum accuracy of only 4/6 even at the higher intervention strength ($\alpha$=125). This variance underscores the value of probing multiple examples and suggests prioritizing heads effective on challenging examples like ’307’ for robust control. Furthermore, as illustrated in Figure 1, intervention strength can impact how much a model is steered towards a specific behaviour.
@@
Figure 2: Sensitivity of specific examples to intervention across attention heads at different intervention strengths. The first row shows results for intervention strength $\alpha=75$ for examples 294 (left), 304 (middle), and 307 (right). The second row shows results for intervention strength $\alpha=125$.
4.3 Hyperparameter screening methods
4.3.1 Layer-wise interventions
For Llama-2, we followed the CAA methodology outlined in (Rimsky et al., 2024), using the same intervention vectors from the paper. We also sweep over all layers intervening one by one with three different intervention strengths and extracting "(A)" or "(B)" to measure the accuracy at each layer, as shown in Figure 3.